As a Physical Education teacher of Hermosa National High School Senior High School, Maam Mary Ann wants to compute for the Body Mass Index of few students as part of her research. As a start, the table below shows the heights of five students in meters. Supposed samples of size 2 are taken from this population of five students, complete the given table and answer the following tasks. Construct the desired tables needed for your solutions and answers.


1
Expert's answer
2021-05-07T11:51:26-0400

With this question we have given a table also,





"\mu = mean\hspace{1mm} of\hspace{1mm} height = \dfrac{7.92}{5} = 1.584"


"\sigma^2 = \sum X^2 - n \times \bar X^2"


"\sigma^2 = 12.59- 5 \times 1.584^2"


"= 0.4472"


Hence, "\sigma = 0.21"